OpenStudy (anonymous):

Kia drew two triangular pyramids in her notebook. The dimensions of the pyramids are shown in the table below.Which statement is true about the pyramids Kia drew?

OpenStudy (anonymous):

They are not similar because the ratio of their corresponding sides is not equal. They are congruent because the ratio of their corresponding sides is not equal. They are similar because the ratio of two of their corresponding sides is the same. They are not congruent because the ratio of their corresponding sides is equal to one.

Directrix (directrix):

They are not similar because the ratio of their corresponding sides is not equal. ---------------- Looking at the sides, form the ratios of what might be corresponding sides by matching the smallest sides of one pyramid's base to those of the other. 12/15 = 16/24 = 20/25 = 10/13 if the figures have a chance of being similar. This is not the case so no similarity or congruence here.

OpenStudy (anonymous):

Thanks